5/16+ 4/5+7/40=
12(4) + (3-6)=
1
Expert's answer
2018-08-30T12:46:09-0400
5/16 + 4/5 + 7/40 = (5*5 + 4*16 + 7*2)/80 = 103/80.
12(4) + (3 - 6) = 48 - 3 = 45.